which prime numbers can make the number 90?
Found 2 solutions by JBarnum, Alan3354:
Answer by JBarnum(2146)

3*3*2*5=90
so 2,3,5

Answer by Alan3354(69443)

90 = 7 + 83
= 11 + 79
= 17 + 73
= 19 + 71
= 23 + 67
= 29 + 61
= 31 + 59
= 37 + 53
= 43 + 47
-----------
And 1 + 89 if you want to count that.
I did positive only.
There's -7 + 97, etc, but prime number problems don't usually include negative numbers.